# Machine Learning with Physics Knowledge for Prediction: A Survey

### v1

@@ 3.4 Operator Learning
where ${\mathsf{F}}$ and ${\mathsf{F}}^{-1}$ denote the Fourier transformation and inverse Fourier transformation, respectively. Like GNOs, a discretized domain representation facilitates an efficient evaluation of equation  (28) by using fast Fourier transformations (FFTs). Given an equidistantly spaced discretization of the domain $\{{\bm{x}}_{k}\}_{k=1}^{N}$, the Fourier coefficients of the discretized Fourier transform of the function ${\bm{v}}_{t-1}$ are ${\hat{\bm{v}}_{t-1}(k)=\sum_{j=0}^{N-1}{\bm{v}}_{t-1}({\bm{x}}_{j+1})e^{-2\pi\mathrm{i}jk/(N)}\in{\mathbb{R}}^{d_{{\bm{v}}_{t-1}}}}$. Likewise, we denote the Fourier coefficients of the kernel by ${\hat{\bm{\kappa}}_{t}(k)\in{\mathbb{R}}^{d_{{\bm{v}}_{t}}\times d_{{\bm{v}}_{t-1}}}}$. Finally, the convolution can be expressed as
${\mathsf{F}}^{-1}({\mathsf{F}}(\kappa_{t})\cdot{\mathsf{F}}({\bm{v}}_{t-1}))({\bm{x}})\approx\textstyle\sum_{k=0}^{N-1}\hat{\bm{\kappa}}(k)\hat{\bm{v}}_{t-1}(k)\,\exp(2\pi\mathrm{i}kn/N).$ (29)
@@
Instead of projecting ${\bm{\kappa}}_{t}$ to the Fourier space, the kernel is directly parameterized in Fourier space. Specifically, the Fourier coefficients of the kernel are parameterized $\hat{\bm{\kappa}}_{{\bm{\theta}}}$. Empirically, it has been found that directly parameterizing the matrix $\hat{\bm{\kappa}}_{{\bm{\theta}}}\in{\mathbb{R}}^{N\times d_{{\bm{v}}_{t}}\times d_{{\bm{v}}_{t-1}}}$ yields the best performance. In addition to their natural frequency domain processing, FNOs demonstrate computational efficiency with a sub-quadratic complexity of $\mathcal{O}(N\log N)$.
@@
However, several studies (Fanaskov and Oseledets, 2022; Bartolucci et al., 2023) emphasize that choosing the training data from a fixed grid representation leads to a systematic bias. In addition, aliasing can occur for FNOs, as the maximum frequency mode that can be distinguished from lower frequency modes is bounded by the grid resolution. Fanaskov and Oseledets (2022) propose spectral neural operators that use Chebychev or trigonometric polynomials.
@@
In recent years, neural operators have gained considerable attention. Different architectures have been proposed based on the recent success in other research disciplines, such as attention-based (Cao, 2021), CNN-based (Raonić et al., 2023), or VAE-based (Seidman et al., 2023) architectures for operator learning. Finally, strides have been made to make neural operators universally applicable in various domains. Liu et al. (2023a) extend FNOs to be agnostic to the domain such that the operator is not restricted to a rectangular domain discretization. Furthermore, neural operator architectures have been extended to be equivariant to rotations, translations, and reflections (Helwig et al., 2023), or explicitly being $SE(3)$-equivariant (Cheng and Peng, 2023). FNOs have also been adopted for general machine learning. FNOs have been incorporated into vision transformers to perform ‘token mixing’ in the Fourier domain rather than in pixel space (Guibas et al., 2022). These adaptive FNOs are developed to be more sparse, use fewer parameters for greater scalability, and have been deployed on weather forecasting tasks (Pathak et al., 2022).
@@
The little regularization and flexibility of neural operators come at the cost of demanding plenty of training data to learn the desired operator. Therefore, one of the ongoing tasks is to incorporate physical priors to facilitate faster and more efficient learning. DeepONets and neural operators have been integrated with physics-informed losses (see section 3.2), thereby enhancing the data-driven operator learning loss with PDE constraints based on the PINN loss ${\mathcal{L}}=\alpha{\mathcal{L}}_{\mathrm{NO}}+(1-\alpha){\mathcal{L}}_{\mathrm{PINN}},\,\alpha\in[0,1]$ (Wang et al., 2021b; Li et al., 2021b). Setting $\alpha=0$ recovers PINNs, but instead of a point-wise mapping, an operator is learned that maps an input function ${\bm{a}}$ to the solution ${\bm{u}}$.
@@
The development of operator learning architectures is strongly influenced by the analysis of linear differential equations. However, the theoretical foundations solely shaped the design of the neural architecture, which is geared towards approximating non-linear operators. In contrast, Koopman operator theory (section 3.5) exclusively addresses dynamical systems represented by linear operators. Gin et al. (2021) demonstrated the close ties between operator learning and Koopman theory by proposing to learn latent representations of input and solution spaces, facilitating the approximation of Green’s function (equation  (24)) for the solution of linear PDEs. For discretizations ${\bm{U}}$ and ${\bm{F}}$ of ${\bm{u}}({\bm{x}})$ and ${\bm{f}}({\bm{x}})$ respectively, they propose learning embeddings ${\bm{u}}_{\bm{\phi}}={\bm{\phi}}_{\bm{\theta}}({\bm{U}})$ and ${\bm{f}}_{\bm{\psi}}={\bm{\psi}}_{\bm{\theta}}({\bm{F}})$ such that ${\bm{L}}{\bm{u}}_{\bm{\phi}}={\bm{f}}_{\bm{\psi}}$. Solving the PDE in this embedding space is simply ${\bm{U}}={\bm{\phi}}_{\bm{\theta}}^{-1}({\bm{L}}^{-1}{\bm{\psi}}_{\bm{\theta}}({\bm{F}}))$, providing ${\bm{\phi}}$ and ${\bm{\psi}}$ can be learned. The subsequent section will discuss the notion of learning effective representations for dynamical systems and the extension of linear dynamical systems theory to nonlinear operators through observable functions.
@@ 3.5 Latent Variable Models for Dynamical Systems
Latent variable models (LVM) (Bishop and Nasrabadi, 2006) refers to the assumption that an observed set of measurements are a consequence of an unobserved latent process. Much of statistics, signal processing and machine learning reduce to problems of this form. In the context of learning PDEs and dynamical systems, there is a desire to reduce complex systems into simpler ones using the idea of a latent coordinate system. This is the motivation behind Koopman operator theory (Koopman, 1931; Koopman and Neumann, 1932), which aims to find a latent coordinate system in which the dynamics become linear, and therefore significantly easier to forecast with. Another motivation is that when learning from measurement data, there is a degree of flexibility in the actual definition of the dynamical system and its internal state, and this ambiguity can be leveraged to learn a dynamical system that may not reflect the physical laws of the process, but may be easier to learn, capture the observed data, or use as a predictive model. In this section, we will discuss learning latent dynamics with both the linear and nonlinear assumption.
@@
Algorithms related to dynamic mode decomposition (DMD) provide entirely data-driven approaches to extract latent coordinate systems from time-series data (Schmid, 2022). Rowley et al. (2009) highlighted the connection between these approaches and Koopman operator theory. Given a temporal sequence $\{{\bm{u}}_{i}\}_{i=1}^{T}$, they operate as supervised learning schemes, addressing least-squares problems of the form
@@
Here, ${\mathcal{J}}({\bm{K}},{\bm{\theta}})$ comprises a one-step prediction loss ${\mathcal{L}}_{\text{Lin}}({\bm{K}},{\bm{\theta}})$ ensuring linearity and a regularisation loss ${\mathcal{L}}_{\text{Reg}}({\bm{K}},{\bm{\theta}})$ scaled by $\alpha\in{\mathbb{R}}^{+}$. The dictionaries ${\bm{\Psi}}^{\prime}:\Omega\rightarrow{\mathbb{C}}^{d_{z}}$ and ${\bm{\Psi}}:\Omega\rightarrow{\mathbb{C}}^{d_{z}}$ represent linear basis models approximating the invariant subspace based on provided feature functions $\{\phi^{j}({\bm{u}}):\Omega\rightarrow{\mathbb{C}}\}_{j=1}^{d_{z}}$. The original DMD framework, introduced by Schmid (2010), considered a one-to-one mapping, i.e. ${\bm{\Psi}}^{\prime}({\bm{u}})={\bm{\Psi}}({\bm{u}})={\bm{u}}$, resulting in linear least-squares problem. Later research, such as extended DMD (EDMD) (Williams et al., 2015) or SINDy (Brunton et al., 2016b), explored more representative features employing non-linear function approximations. However, choosing ${\bm{\Psi}}^{\prime}$ and ${\bm{\Psi}}$ is challenging, commonly known as the bias-variance problem in machine learning (Bishop and Nasrabadi, 2006). It results in conflicting goals between (i) extracting a valid KIS and (ii) not overfitting to the given data (Otto and Rowley, 2019). Therefore, Li et al. (2017) and Yeung et al. (2019) proposed dictionaries with a partly trainable set of features, i.e., $\{\phi^{j}_{\bm{\theta}}({\bm{u}}):\Omega\rightarrow{\mathbb{C}}\}_{j=1}^{k}$ and $k<d_{z}$. EDMD with dictionary learning (EDMD-DL) considers shallow networks as function approximations and utilizes an $L^{2}$-regularization for ${\bm{\theta}}$ (Li et al., 2017). In contrast, Yeung et al. (2019) proposed deep-DMD employing deep learning architectures and a $L^{1}$-regularization. Botch methods iterate between a ridge regression to estimate ${\bm{K}}$ and a nonlinear gradient descent to optimize ${\bm{\theta}}$. Lastly, Terao et al. (2021) introduced a variant of EDMD-DL, replacing the MLPs with NODEs (section 3.1). These approaches have demonstrated promising results in capturing chaotic behavior, such as that exhibited by the Duffing oscillator (Williams et al., 2015). Also, they found practical applications in fluid mechanics, for instance, in systems described by the Kuramoto-Sivashinsky equation (Li et al., 2017). However, they share the assumption of a full state observable i.e. ${\bm{g}}({\bm{u}})={\bm{u}}$. This assumption is implemented in two ways. Firstly, through a linear reconstruction using the estimated eigenfunctions and Koopman modes. Alternatively, by incorporating a constant feature $\phi({\bm{u}})={\bm{u}}$ into the dictionary.
@@
In recent years, researchers explored the application of autoencoder architectures to extract a concise yet informative subspace from time series data (Takeishi et al., 2017; Lusch et al., 2018; Alford-Lago et al., 2022). As a result, they relaxed the assumption of a full state observable. The linear reconstruction becomes obsolete and a non-linear reconstruction takes place using the decoder network. The objective from equation  (32) extends to
@@
with ${\mathcal{L}}_{\text{Recon}}({\bm{\theta}}_{e},{\bm{\theta}}_{d})$ and ${\mathcal{L}}_{\text{Pred}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ being a reconstruction loss and a prediction loss, respectively. The former penalizes inaccurate reconstructions of the system state, i.e. ${{\bm{u}}_{i}\approx(d_{{\bm{\theta}}_{d}}\circ e_{{\bm{\theta}}_{e}}){\bm{u}}_{i}}$. The latter aims to correctly predict the next state after propagating the current one through the architecture, i.e. ${{\bm{u}}_{i+1}\approx(d_{{\bm{\theta}}_{d}}\circ{\bm{K}}\circ e_{{\bm{\theta}}_{e}}){\bm{u}}_{i}}$. Several works, including Takeishi et al. (2017); Lusch et al. (2018) and Alford-Lago et al. (2022) focused on forward prediction losses ${\mathcal{L}}_{\text{Pred}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ and ${\mathcal{L}}_{\text{Lin}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$. Azencot et al. (2020) studied the effect of an additional backward prediction loss, i.e. ${\bm{u}}_{i-1}\approx(d_{{\bm{\theta}}_{d}}\circ{\bm{D}}\circ e_{{\bm{\theta}}_{e}}){\bm{u}}_{i}$. They introduced an additional operator for the backward dynamics ${\bm{D}}$ satisfying ${\bm{K}}\circ{\bm{D}}={\bm{D}}\circ{\bm{K}}=\mathbb{I}$. However, these works solely looked at one-step predictions. Otto and Rowley (2019), on the other hand, introduced an autoencoder featuring linear recurrent latent embeddings. The proposed linearly-recurrent autoencoder network (LRAN) enhances the recurrent losses ${\mathcal{L}}_{\text{Pred}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ and ${\mathcal{L}}_{\text{Lin}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ in equation  (33) by considering a weighted average over $\tau$-step predictions. Similarly, Wang et al. (2022c) and Liu et al. (2024) both used recurrent structures to learn KIS and operators. However, they employ a composition of latent coordinate systems, each designed to capture different aspects of the time series data. The Koopman neural forecaster (KNF) decomposes the embedding to capture local and global behavior patterns separately (Wang et al., 2022c). In contrast, Liu et al. (2024) proposed a hierarchy of Koopman forecaster blocks (‘Koopa’). In each block, a Fourier filter separates the input signal to capture time-varying and time-invariant components. While these methods were effective for applications like weather forecasting (Liu et al., 2024) and fluid mechanics, e.g., in modeling wake flow around cylinders (Otto and Rowley, 2019), they showed efficacy in quantifying the impacts of uncertainties in both model construction and forecasting. Therefore, several works, including Morton et al. (2019) and Pan and Duraisamy (2020), proposed using variational autoencoder (VAE) architectures (Kingma and Welling, 2014). These approaches enable the learning of a generative model that closely resembles the distribution over dynamics in the latent space.
@@
Eigenvalues are crucial for understanding the temporal behavior of the dynamics in the latent embedding. Previous works, including Liu et al. (2024) or Otto and Rowley (2019), have not explicitly addressed stability. Consequently, the latent dynamics were only approximately linear and prone to instability. Therefore, recent research proposes physics-informed architectures to promote properties like stability. Erichson et al. (2019) suggested an autoencoder framework targeting stable latent Koopman dynamics, following Lyapunov stability principles. Pan and Duraisamy (2020), on the other hand, employed a skew-symmetric representation of the Koopman operator. This approach inherently forms stable latent dynamics and deviates from Erichson et al. (2019), where stability is integrated as a soft constraint within the objective function equation  (33). For quasi-periodic systems, works like Lange et al. (2021) express the objective in the frequency domain leveraging the fast Fourier transformation (FFT). This dual formulation achieves globally linear latent embeddings by construction. To incorporate further physical principles such as conservation or measure preservation, Baddoo et al. (2023) or Colbrook (2023) reframed the regression problem from equation  (32) as Procrustes problem. This formulation ensures that the solution lies on a manifold spanned by the physically imposed constraints. While demonstrating convergence guarantees and promising results for systems such as the Volterra integro-differential equation or linearized Navier–Stokes equations, these approaches rely on the choice of feature functions and assume linear reconstruction.

### v2

@@ 3.4 Operator Learning
where ${\mathsf{F}}$ and ${\mathsf{F}}^{-1}$ denote the Fourier transformation and inverse Fourier transformation, respectively Like GNOs, a discretized domain representation facilitates an efficient evaluation of equation (28) by using fast Fourier transformations (FFTs). Given an equidistantly spaced discretization of the domain $\{{\bm{x}}_{k}\}_{k=1}^{N}$, the Fourier coefficients of the discretized Fourier transform of the function ${\bm{v}}_{t-1}$ are ${\hat{\bm{v}}_{t-1}(k)=\sum_{j=0}^{N-1}{\bm{v}}_{t-1}({\bm{x}}_{j+1})e^{-2\pi\mathrm{i}jk/(N)}\in{\mathbb{R}}^{d_{{\bm{v}}_{t-1}}}}$. Likewise, we denote the Fourier coefficients of the kernel by ${\hat{\bm{\kappa}}_{t}(k)\in{\mathbb{R}}^{d_{{\bm{v}}_{t}}\times d_{{\bm{v}}_{t-1}}}}$. Finally, the convolution can be expressed as
${\mathsf{F}}^{-1}({\mathsf{F}}(\kappa_{t})\cdot{\mathsf{F}}({\bm{v}}_{t-1}))({\bm{x}})\approx\textstyle\sum_{k=0}^{N-1}\hat{\bm{\kappa}}_{t}(k)\hat{\bm{v}}_{t-1}(k)\,\exp(2\pi\mathrm{i}kn/N).$ (29)
@@
Instead of projecting $\kappa_{t}$ to the Fourier space, the kernel is directly parameterized in Fourier space. Specifically, the Fourier coefficients of the kernel are parameterized $\hat{\bm{\kappa}}_{{\bm{\theta}}}$. Empirically, it has been found that directly parameterizing the matrix $\hat{\bm{\kappa}}_{{\bm{\theta}}}\in{\mathbb{R}}^{N\times d_{{\bm{v}}_{t}}\times d_{{\bm{v}}_{t-1}}}$ yields the best performance. In addition to their natural frequency domain processing, FNOs demonstrate computational efficiency with a sub-quadratic complexity of $\mathcal{O}(N\log N)$.
@@
However, several studies (Fanaskov and Oseledets, 2023; Bartolucci et al., 2023) emphasize that choosing the training data from a fixed grid representation leads to a systematic bias. In addition, aliasing can occur for FNOs, as the maximum frequency mode that can be distinguished from lower frequency modes is bounded by the grid resolution. Fanaskov and Oseledets (2023) propose spectral neural operators that use Chebyshev or trigonometric polynomials. Additionally, multi-wavelet operators (Gupta et al., 2021) extend FNOs by utilizing multi-resolution wavelets to represent the kernel operator with a piecewise polynomial basis. This results in a sparse kernel representation, improving the operator learning performance.
@@
In recent years, neural operators have gained considerable attention. Different architectures have been proposed based on the recent success in other research disciplines, such as attention-based (Cao, 2021), CNN-based (Raonić et al., 2023), or VAE-based (Seidman et al., 2023) architectures for operator learning. Finally, strides have been made to make neural operators universally applicable in various domains compliant to specific pre-conditions. As FNOs operate on uniform grids, domain-agnostic methods have been proposed that either map complex topologies to a latent uniform grid for the NO (Li et al., 2023a; b) or apply the NO on an extended uniform grid and masking out points outside the domain (Liu et al., 2023a). Furthermore, as neural operators are fully data-driven approaches, there are no guarantees that the learned solution operator adheres to the boundary value conditions of the considered PDE. Therefore, Saad et al. (2023) show that specific parameterizations of the kernel operator 25 can ensure compliance with Dirichlet, Neumann, and periodic boundary conditions while Négiar et al. (2023) propose to employ a differentiable final layer which guarantees that the boundary constraints are not violated. Further geometric biases have been employed with neural operator architectures to be equivariant to rotations, translations, and reflections (Helwig et al., 2023), or explicitly being $\mathrm{SE}(3)$-equivariant (Cheng and Peng, 2023). FNOs have also been adopted for general machine learning. FNOs have been incorporated into vision transformers to perform ‘token mixing’ in the Fourier domain rather than in pixel space (Guibas et al., 2022). These adaptive FNOs are developed to be more sparse, use fewer parameters for greater scalability, and have been deployed on weather forecasting tasks (Pathak et al., 2022).
@@
The little regularization required for training and high flexibility of neural operators come at the cost of demanding plenty of training data to learn the desired operator. Therefore, one of the ongoing tasks is to incorporate physical priors to facilitate faster and more efficient learning. DeepONets and neural operators have been integrated with physics-informed losses, see section 3.3, thereby enhancing the data-driven operator learning loss with PDE constraints based on the PINN loss ${\mathcal{L}}=\alpha{\mathcal{L}}_{\mathrm{NO}}+(1-\alpha){\mathcal{L}}_{\mathrm{PINN}},\,\alpha\in[0,1]$ (Wang et al., 2021c; Li et al., 2021b). Setting $\alpha=0$ recovers PINNs, but instead of a point-wise mapping, an operator is learned that maps an input function ${\bm{a}}$ to the solution ${\bm{u}}$. However, Saad et al. (2023) show that the auxiliary PINN loss does not always benefit the training of neural operators and may hinder its final performance. Du et al. (2024) highlight the shortcomings of the PINN loss as it requires possibly multiple differentiations through the network architecture. Thus, the authors propose to replace the PINN loss term with a new loss based on spectral theory. The proposed loss evaluates the residual term of the PINN loss in its spectral representations, which can be shown to simplify gradient propagation through the network and thus improve the optimization of neural operators with added physics-inspired losses.
The study of linear differential equations, particularly through Green’s functions, has significantly shaped the development of operator learning architectures. Here, nonlinear operator approximation is achieved by compositions of linear operators with nonlinear activations. In the next section (section 3.5), we introduce a related but distinct approach, Koopman theory, which also leverages linear operators to approximate nonlinear dynamical systems. Unlike the compositional method, Koopman theory seeks to embed the domain into a latent space where the dynamics are governed by a linear, though potentially infinite-dimensional, operator.
@@
Gin et al. (2021) demonstrated the close ties between operator learning and Koopman theory by proposing to learn latent representations of input and solution spaces, facilitating the approximation of Green’s function (equation (24)) for the solution of linear PDEs. For discretizations ${\bm{U}}$ and ${\bm{F}}$ of ${\bm{u}}({\bm{x}})$ and ${\bm{f}}({\bm{x}})$ respectively, they propose learning embeddings ${\bm{u}}_{\bm{\phi}}={\bm{\phi}}_{\bm{\theta}}({\bm{U}})$ and ${\bm{f}}_{\bm{\psi}}={\bm{\psi}}_{\bm{\theta}}({\bm{F}})$ such that ${\bm{L}}{\bm{u}}_{\bm{\phi}}={\bm{f}}_{\bm{\psi}}$. Solving the PDE in this embedding space is simply ${\bm{U}}={\bm{\phi}}_{\bm{\theta}}^{-1}({\bm{L}}^{-1}{\bm{\psi}}_{\bm{\theta}}({\bm{F}}))$, providing ${\bm{\phi}}$ and ${\bm{\psi}}$ can be learned. The subsequent section will discuss the notion of learning effective representations for dynamical systems and the extension of linear dynamical systems theory to nonlinear operators through observable functions.
@@ 3.5 Latent Variable Models for Dynamical Systems
Latent variable models (LVM) (Bishop and Nasrabadi, 2006) refers to the assumption that an observed set of measurements are a consequence of an unobserved latent process. Much of statistics, signal processing and machine learning reduce to problems of this form. In the context of learning PDEs and dynamical systems, there is a desire to reduce complex systems into simpler ones using the idea of a latent coordinate system. Thus, LVMs motivate a different view on previously discussed methods for operator learning on finite-dimensional embeddings, see section 3.4, and physically-structured latent coordinate systems, see section 3.2. These methods construct latent coordinate representations enhancing the interpretability of system dynamics and enabling more efficient computational analysis. Building on this motivation, the following section surveys Koopman operator theory, which aims to find a latent coordinate system in which the dynamics become linear (Koopman, 1931; Koopman and Neumann, 1932). The linearization simplifies the forecasting and enables the use of well-established techniques from linear algebra for analyzing Another motivation is that when learning from measurement data, there is a degree of flexibility in the actual definition of the dynamical system and its internal state, and this ambiguity can be leveraged to learn a dynamical system that may not reflect the physical laws of the process, but may be easier to learn, capture the observed data, or use as a predictive model. In this section, we will discuss learning latent dynamics with both the linear and nonlinear assumption.
@@
Algorithms related to dynamic mode decomposition (DMD) provide entirely data-driven approaches to extract latent coordinate systems from time-series data (Schmid, 2022). They decompose complex systems into spatio-temporal coherent structures employing singular value decomposition (Kutz et al., 2016). Rowley et al. (2009) highlighted the connection between these approaches and Koopman operator theory. Given a temporal sequence $\{{\bm{u}}_{i}\}_{i=1}^{T}$, they operate as supervised learning schemes, addressing least-squares problems of the form
@@
Here, ${\mathcal{J}}({\bm{K}},{\bm{\theta}})$ comprises a one-step prediction loss ${\mathcal{L}}_{\text{Lin}}({\bm{K}},{\bm{\theta}})$ ensuring linearity and a regularization loss ${\mathcal{L}}_{\text{Reg}}({\bm{K}},{\bm{\theta}})$ scaled by $\alpha\in{\mathbb{R}}^{+}$. The dictionaries ${\bm{\Psi}}^{\prime}:\Omega\rightarrow{\mathbb{C}}^{d_{z}}$ and ${\bm{\Psi}}:\Omega\rightarrow{\mathbb{C}}^{d_{z}}$ represent linear basis models approximating the invariant subspace based on provided feature functions $\{\phi^{j}({\bm{u}}):\Omega\rightarrow{\mathbb{C}}\}_{j=1}^{d_{z}}$. The original DMD framework, introduced by Schmid (2010), considered a one-to-one mapping, i.e. ${\bm{\Psi}}^{\prime}({\bm{u}})={\bm{\Psi}}({\bm{u}})={\bm{u}}$, resulting in linear least-squares problem. Later research, such as extended DMD (EDMD) (Williams et al., 2015) or SINDy (Brunton et al., 2016b), explored more representative features employing nonlinear function approximations. However, choosing ${\bm{\Psi}}^{\prime}$ and ${\bm{\Psi}}$ is challenging, commonly known as the bias-variance problem in machine learning (Bishop and Nasrabadi, 2006). It results in conflicting goals between (i) extracting a valid KIS and (ii) not overfitting to the given data (Otto and Rowley, 2019). Therefore, Li et al. (2017) and Yeung et al. (2019) proposed dictionaries with a partly trainable set of features, i.e., $\{\phi^{j}_{\bm{\theta}}({\bm{u}}):\Omega\rightarrow{\mathbb{C}}\}_{j=1}^{k}$ and $k<d_{z}$. EDMD with dictionary learning (EDMD-DL) considers shallow networks as function approximations and utilizes an $L^{2}$-regularization for ${\bm{\theta}}$ (Li et al., 2017). In contrast, Yeung et al. (2019) proposed deep-DMD employing deep learning architectures and a $L^{1}$-regularization. Batch methods iterate between a ridge regression to estimate ${\bm{K}}$ and a nonlinear gradient descent to optimize ${\bm{\theta}}$. Furthermore, Terao et al. (2021) introduced a variant of EDMD-DL, replacing the MLPs with NODEs (section 3.1). These approaches have demonstrated promising results in capturing chaotic behavior, such as that exhibited by the Duffing oscillator (Williams et al., 2015). Also, they found practical applications in fluid mechanics, for instance, in systems described by the Kuramoto-Sivashinsky equation (Li et al., 2017). However, they share the assumption of a full state observable i.e. ${\bm{g}}({\bm{u}})={\bm{u}}$. This assumption is implemented in two ways. Firstly, through a linear reconstruction using the estimated eigenfunctions and Koopman modes. Alternatively, by incorporating a constant feature $\phi({\bm{u}})={\bm{u}}$ into the dictionary. However, a significant challenge remains in the robust forecasting of out-of-distribution dynamics. Mouli et al. (2024a) addressed this problem employing a meta-learning framework that learns relevant feature functions across tasks. Similarly to SINDy (Brunton et al., 2016b), their approach employs nonlinear function approximators together with task-specific coefficient matrices to learn the dynamics for each task. In addition, they introduce a task-independent indicator matrix that learns the global causal structure across all tasks. Given the causal information, efficient adaptation to new tasks is achieved by learning only the task-specific coefficients. This method demonstrates the potential of meta-learning, discussed in section 4, to address challenges such as out-of-distribution dynamics.
@@
In recent years, researchers explored the application of autoencoder architectures to extract a concise yet informative subspace from time series data (Takeishi et al., 2017; Lusch et al., 2018; Alford-Lago et al., 2022). As a result, they relaxed the assumption of a full state observable. The linear reconstruction becomes obsolete and a nonlinear reconstruction takes place using the decoder network. The objective from equation (32) extends to
@@
with ${\mathcal{L}}_{\text{Recon}}({\bm{\theta}}_{e},{\bm{\theta}}_{d})$ and ${\mathcal{L}}_{\text{Pred}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ being a reconstruction loss and a prediction loss, respectively. The former penalizes inaccurate reconstructions of the system state, i.e. ${{\bm{u}}_{i}\approx(d_{{\bm{\theta}}_{d}}\circ e_{{\bm{\theta}}_{e}}){\bm{u}}_{i}}$. The latter aims to correctly predict the next state after propagating the current one through the architecture, i.e. ${{\bm{u}}_{i+1}\approx(d_{{\bm{\theta}}_{d}}\circ{\bm{K}}\circ e_{{\bm{\theta}}_{e}}){\bm{u}}_{i}}$. Several works, including Takeishi et al. (2017); Lusch et al. (2018) and Alford-Lago et al. (2022) focused on forward prediction losses ${\mathcal{L}}_{\text{Pred}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ and ${\mathcal{L}}_{\text{Lin}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$. Azencot et al. (2020) studied the effect of an additional backward prediction loss, i.e. ${\bm{u}}_{i-1}\approx(d_{{\bm{\theta}}_{d}}\circ{\bm{D}}\circ e_{{\bm{\theta}}_{e}}){\bm{u}}_{i}$. They introduced an additional operator for the backward dynamics ${\bm{D}}$ satisfying ${\bm{K}}\circ{\bm{D}}={\bm{D}}\circ{\bm{K}}=\mathbb{I}$. However, these works solely looked at one-step predictions. Otto and Rowley (2019), on the other hand, introduced an autoencoder featuring linear recurrent latent embeddings. The proposed linearly-recurrent autoencoder network (LRAN) enhances the recurrent losses ${\mathcal{L}}_{\text{Pred}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ and ${\mathcal{L}}_{\text{Lin}}({\bm{\theta}}_{e},{\bm{\theta}}_{d},{\bm{K}})$ in equation (33) by considering a weighted average over $\tau$-step predictions. Similarly, Wang et al. (2022c) and Liu et al. (2024a) both used recurrent structures to learn KIS and operators. However, they employ a composition of latent coordinate systems, each designed to capture different aspects of the time series data. The Koopman neural forecaster (KNF) decomposes the embedding to capture local and global behavior patterns separately (Wang et al., 2022c). In contrast, Liu et al. (2024a) proposed a hierarchy of Koopman forecaster blocks (‘Koopa’). In each block, a Fourier filter separates the input signal to capture time-varying and time-invariant components. While these methods were effective for applications like weather forecasting (Liu et al., 2024a) and fluid mechanics, e.g., in modeling wake flow around cylinders (Otto and Rowley, 2019), they showed efficacy in quantifying the impacts of uncertainties in both model construction and forecasting. Therefore, several works, including Morton et al. (2019) and Pan and Duraisamy (2020), proposed using variational autoencoder (VAE) architectures (Kingma and Welling, 2014). These approaches enable the learning of a generative model that closely resembles the distribution over dynamics in the latent space.
@@
Eigenvalues are crucial for understanding the temporal behavior of the dynamics in the latent embedding. Previous works, including Liu et al. (2024a) or Otto and Rowley (2019), have not explicitly addressed stability. As a consequence, the locally-linear latent dynamics were prone to instability. Therefore, recent research proposes physics-informed architectures to promote properties like stability. Erichson et al. (2019) suggested an autoencoder framework targeting stable latent Koopman dynamics, following Lyapunov stability principles. Pan and Duraisamy (2020), on the other hand, employed a skew-symmetric representation of the Koopman operator. This approach inherently forms stable latent dynamics and deviates from Erichson et al. (2019), where stability is integrated as a soft constraint within the objective function equation (33). For quasi-periodic systems, works like Lange et al. (2021) express the objective in the frequency domain leveraging the fast Fourier transformation (FFT). This dual formulation achieves globally linear latent embeddings by construction. To incorporate further physical principles such as conservation or measure preservation, Baddoo et al. (2023) or Colbrook (2023) reframed the regression problem from equation (32) as Procrustes problem. This formulation ensures that the solution lies on a manifold spanned by the physically imposed constraints. While demonstrating convergence guarantees and promising results for systems such as the Volterra integro-differential equation or linearized Navier–Stokes equations, these approaches rely on the choice of feature functions and assume linear reconstruction.